A sample size of 400 is taken from a population x with a u=4 and o=2.Find the probability that the sample mean is less than 3.85

1
Expert's answer
2022-05-02T15:27:10-0400
"P(\bar{X}<3.85)=P(Z<\dfrac{3.85-4}{2/\sqrt{400}})"

"=P(Z<-1.5)\approx0.066807"